# Boundary-driven quantum systems near the Zeno limit: steady states and long-time behavior

### v1

@@ 6 Davies’ oscillation averaging operation
As in [13], for any super-operator $\mathcal{T}$ on $\widehat{\mathcal{H}}_{B}$, define $\mathcal{T}^{\sharp}$ by
@@
We are mainly interested in $\mathcal{D}_{P}^{\sharp}$ where $\mathcal{D}_{P}$ is a projected Lindbladian generator on $\widehat{\mathcal{H}}_{B}$. However, many properties of Davies’ coherent averaging operation are valid for all super-operator $\mathcal{T}$ on $\widehat{\mathcal{H}}_{B}$.
@@
Let $\mathscr{D}R=\sum_{\ell=1}^{r}\left(2V_{\ell}RV_{\ell}^{\dagger}-V_{\ell}^{\dagger}V_{\ell}R-RV_{\ell}^{\dagger}V_{\ell}\right)-i[H_{L},R]$} be any Lindblad generator on $\widehat{\mathcal{H}}_{B}$. Let $\mathscr{D}^{\sharp}$ be defined in terms of $\mathcal{D}$ through (98). Let $H_{P}=\sum_{\lambda\in\sigma(H_{P})}\lambda P_{\lambda}$ be the spectral decomposition of $H_{P}$ so that $P_{\lambda}$ is the orthogonal projection on the the eigenspace with eigenvalue $\lambda$, and $\sigma(H_{p})$ is the spectrum of $H_{P}$.
@@
If $P_{\lambda+\omega}\neq 0$, $\omega=\mu-\lambda$ for some $\mu\in\sigma(H_{P})$. For all $\lambda,\mu\in\sigma(H_{P})$, there is exactly one $\omega\in\sigma(i\mathcal{K})$ such that $\lambda+\omega=\mu$. To see this, note that $\omega_{0}:=\mu-\lambda\in\sigma(i\mathcal{K})$, so that there is at least one such $\omega$. But every other $\omega\in\sigma(i\mathcal{K})$ satisfies either $\omega>\omega_{0}$ or $\omega<\omega_{0}$. Therefore, for every other $\omega$, $\lambda+\omega>\mu$ or $\lambda+\omega<\mu$, and hence
@@
We consider a very simple 2 qubit system. Let $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ both be $\mathbbm{C}^{2}$ with the usual inner product. Let $\small|0\rangle:=\left(\begin{array}[]{c}0\cr 1\end{array}\right)$} and $\small|1\rangle:=\left(\begin{array}[]{c}1\cr 0\end{array}\right)$}. Let $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ be the Pauli matrices, and define
@@ 8 Application to the computation of steady states
$P_{\mathcal{W}}=\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}e^{t\mathcal{K}_{P}}{\rm d}t\ .$ (132)
It follows that $P_{\mathcal{W}}$ is a CPTP operation, and hence $\|P_{\mathcal{W}}\|_{1\ to1}=1$.
@@
$\textnormal{Tr}_{A}[\mathcal{K}\bar{n}_{0}]=\textnormal{Tr}_{A}[-\mathcal{K}\mathcal{S}\mathcal{K}(\pi\otimes\bar{R})]+\textnormal{Tr}_{A}[\mathcal{K}\pi_{A}\otimes X_{0}]=\mathcal{D}_{P}\bar{R}+\mathcal{K}_{P}X_{0}=\mathcal{D}_{P}\bar{R}+\mathcal{K}_{P}V_{0}\ .$ (139)
Therefore, there is no constraint at this point on $W_{0}$, but $V_{0}$ must be chosen so that $\mathcal{K}_{P}V_{0}=-\mathcal{D}_{P}\bar{R}$. By part (1) of Lemma 8.1, since $\bar{R}\in{\rm ker}(\mathcal{K}_{P})\cap{\rm ker}(\mathcal{D}_{P}^{\sharp})$, $\mathcal{D}_{P}(\bar{R})\in{\rm ran}(\mathcal{K}_{P})$, and then by part (2) of Lemma 8.1, there is a unique $V_{0}\in\mathcal{V}$ such that $\mathcal{K}V_{0}=-\mathcal{D}_{P}\bar{R}$, and $\|V_{0}\|_{1}\leqslant b^{-1}\|\mathcal{D}_{P}\bar{R}\|_{1}$, from which (137) follows. ∎
Now let $\bar{n}_{0}=\tilde{m}_{0}+\pi_{A}\otimes W_{0}$ as in Lemma 8.2, so that $\mathcal{D}\bar{n}_{1}=-\mathcal{K}\bar{n}_{0}$ is solvable. The general solution has the form
@@
Defining $\tilde{m}_{1}:=-\mathcal{S}\mathcal{K}\bar{n}_{0}+\pi_{A}\otimes V_{1}=-\mathcal{S}\mathcal{K}\tilde{m}_{0}+\pi_{A}\otimes W_{0}+\pi_{A}\otimes V_{1}$}, for any choice of $W_{1}\in\mathcal{W}$, $\mathcal{D}\bar{n}_{2}=-\mathcal{K}(\tilde{m}_{1}+W_{1})$ is solvable, and
@@
For $1\leqslant\ell\leqslant k$, define
$\tilde{m}_{\ell}:=-\mathcal{S}\mathcal{K}n_{\ell-1}+\pi_{A}\otimes V_{\ell}\ ,$ (145)
so that $\bar{n}_{\ell}=\tilde{m}_{\ell}+\pi_{A}\otimes W_{\ell}$, and let $\tilde{m}_{0}$ be given as in Lemma 8.2. Suppose that for all $1\leqslant\ell\leqslant k$,
@@
Then since $\|\bar{n}_{k}\|_{1}\leqslant\|\tilde{m}_{k}\|_{1}+\|W_{k}\|_{1}$.
@@
From (110), it follows that $\mathcal{D}_{P}(\pi_{A})=0$, and
$\mathcal{D}_{P}(\sigma_{-})=-\sigma_{-}\ ,\quad\mathcal{D}_{P}\sigma_{+}=-\sigma_{+}\quad{\rm and}\quad\mathcal{D}_{P}(\sigma_{3})=-2(\sigma_{3})\ .$ (149)
Therefore, $\mathcal{D}_{P}$ is ergodic with unique steady state $\pi_{A}$, and its spectral gap is $1$.
The projected Lindblad generator is $\mathcal{K}_{P}+\gamma^{-1}\mathcal{D}_{P}$ which in the $\{\mathds{1},\sigma_{1},\sigma_{2},\sigma_{3}\}$ basis, and with $t$ denoting $\tanh(\beta/2)$, has the matrix representation
@@
Note also that the real parts of the non-zero eigenvalues of $\gamma\mathcal{L}_{P,\gamma}$ are 1-1 and $\tfrac{3}{2}$, which are exactly the non-zero eigenvalues of $\mathcal{D}_{P}^{\sharp}$. This is consistent with, but does not imply, the relation between the mixing times in Theorem 7.5.
@@ Appendix C Generalized inverses of Lindblad generators
$e^{t\mathcal{D}_{A}}=\sum_{j=0}^{n_{A}}e^{t\lambda_{j}}|Y_{j}\rangle\langle X_{j}|\ ,$ (164)
@@ Appendix D Volterra integral operators and the comparison of evolution equations
$\|\mathscr{H}\|_{\mathscr{C}\to\mathscr{C}}=\sup\{|\!|\!|\mathscr{H}A|\!|\!|\ :\ |\!|\!|A|\!|\!|\leqslant 1\}\ .$ (174)
@@
$\|\mathscr{H}-\widetilde{\mathscr{H}}\|_{\mathscr{C}\to\mathscr{C}}\leqslant\sup_{0\leqslant\tau\leqslant T}\left\{\ \int_{0}^{\tau}\|\mathcal{G}(\tau,\sigma)-\widetilde{\mathcal{G}}(\tau,\sigma)\|_{\mathcal{X}\to\mathcal{X}}{\rm d}\sigma\ \right\}\ .$ (177)
@@ Appendix E Mixing times
Let $\mathcal{L}$ be a Lindblad generator acting on $\widehat{\mathcal{H}}$, $\mathcal{H}$ a finite dimensional Hilbert space. By by (173), if $\mathcal{L}$ is ergodic and gapped with unique steady state $\pi$, $t_{{\rm mix}}(\mathcal{L})<\infty$ . The next lemma provides a converse.
@@
Let $\mathcal{L}$ be a Lindblad generator acting on $\widehat{\mathcal{H}}$. Assume only that $t_{{\rm mix}}(\mathcal{L})<\infty$. Then $\mathcal{L}$ has a unique steady state $\pi$, and ${\rm ker}(\mathcal{L})$ is spanned by $\pi$. Moreover, $\mathcal{L}$ has no purely imaginary eigenvalues so that in our finite dimensional setting $\mathcal{L}$ is ergodic and gapped. In particular, if $\mathcal{L}\rho-0$ and $\textnormal{Tr}[\rho]=1$, then $\rho=\pi$.
@@
Every Lindblad generator acting on $\widehat{\mathcal{H}}$ has at least one steady state, and since $t_{{\rm mix}}(\mathcal{L})<\infty$, there is a unique steady state $\pi$. Since for all $X$, $\mathcal{L}(X^{\dagger})=\mathcal{L}(X)^{\dagger}$, if $X\in{\rm ker}(\mathcal{L})$ then $X^{\dagger}\in{\rm ker}(\mathcal{L})$. Hence if $\pi$ does not span ${\rm ker}(\mathcal{L})$ there would be a self-adjoint $X\in{\rm ker}(\mathcal{L})$ that is linearly independent of $\pi$. Subtracting an appropriate multiple of $\pi$ from $X$ and normalizing, we may assume that $\textnormal{Tr}[X]=0$ but $\|X\|_{1}=2$. Let $X=X_{+}-X_{-}$ be the spectral decomposition of $X$ into its positive and negative parts. Then $X_{+}$ and $X_{-}$ are both density matrices. Since $\mathcal{L}X=0$, for all $t>0$,
@@
Therefore, $2=\|X\|_{1}=2d_{{\rm TV}}(e^{t\mathcal{L}}X_{+},e^{t\mathcal{L}}X_{-})$, but for $t>t_{{\rm mix}}(\mathcal{L})$, $d_{{\rm TV}}(e^{t\mathcal{L}}X_{+},e^{t\mathcal{L}}X_{-})<\tfrac{1}{4}$. This contradiction proves that $\pi$ span ${\rm ker}(\mathcal{L})$.
@@
We next show that when $t_{{\rm mix}}(\mathcal{L})<\infty$, so that there is a unique stationary state $\pi$, $d_{{\rm TV}}(e^{kt_{{\rm mix}}(\mathcal{L})})$ decreases to zero exponentially fast in $k$. For $t>0$, let $P_{t}:=e^{t\mathcal{L}}$. In preparation for this, define the super-operator $\Pi$ by $\Pi X:=\textnormal{Tr}[X]\pi$. Then $\Pi$ is a CPTP map. It easy to write down an explicit Kraus representation showing this, but also $\Pi=\lim_{t\to\infty}P_{t}$, and each $P_{t}$ is CPTP. A super-operator $\mathcal{T}$ is said to be Hermitian in case $\mathcal{T}(X)^{\dagger}=\mathcal{T}(X^{\dagger})$ for all $X\in\widehat{\mathcal{H}}$. Since all CPTP operations are Hermitian (see e.g. [9, Lemma 5.2]), and hence $P_{t}-\Pi$ is Hermitian.
@@
Let $\rho_{0},\rho_{1}$ be any two density matrices on $\mathcal{H}$. For all $0<\epsilon<\tfrac{1}{2}$, and all $t\geqslant t_{{\rm mix}}(\mathcal{L})$, and all positive integers $k$,
@@
Now suppose that $t>t_{{\rm mix}}(\mathcal{L},\epsilon)$. By the definition of $t_{{\rm mix}}(\mathcal{L})$, for any density matrix $\rho$,
@@
$t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ whenever $\mathcal{L}$ is ergofic and gapped. Moreover, by Theorem E.2 the definition of mixing time used here differs from the one defined in terms of (187) by at most a factor of $2$. Again, our applications reuire a formulation that does not refer to a putative unique invariant measure. Many variants of the mixing time are discussed in the classical literature on Markov chains, and the classical analog of the one considered here occurs in the foundational paper [2] of Aldous

### v2

@@ 6 Davies’ oscillation averaging operation
As in [13], for any super-operator $\mathcal{T}$ on $\mathcal{H}_{B}$, define $\mathcal{T}^{\sharp}$ by
@@
We are mainly interested in $\mathcal{D}_{P}^{\sharp}$ where $\mathcal{D}_{P}$ is a projected Lindbladian generator on $\widehat{\mathcal{H}}_{B}$. However, many properties of Davies’ coherent averaging operation are valid for all super-operators $\mathcal{T}$ on $\mathcal{H}_{B}$.
@@
Let $\mathscr{D}R=\sum_{\ell=1}^{r}\left(2V_{\ell}RV_{\ell}^{\dagger}-V_{\ell}^{\dagger}V_{\ell}R-RV_{\ell}^{\dagger}V_{\ell}\right)-i[H_{L},R]$} be any Lindblad generator on $\widehat{\mathcal{H}}_{B}$. Let $\mathscr{D}^{\sharp}$ be defined in terms of $\mathscr{D}$ through (93). Let $H_{P}=\sum_{\lambda\in\sigma(H_{P})}\lambda P_{\lambda}$ be the spectral decomposition of $H_{P}$ so that $P_{\lambda}$ is the orthogonal projection on the the eigenspace with eigenvalue $\lambda$, and $\sigma(H_{P})$ is the spectrum of $H_{P}$.
@@
If $P_{\lambda+\omega}\neq 0$, $\omega=\mu-\lambda$ for some $\mu\in\sigma(H_{P})$. For all $\lambda,\mu\in\sigma(H_{P})$, there is exactly one $\omega\in\sigma(i\mathcal{K}_{P})$ such that $\lambda+\omega=\mu$. To see this, note that $\omega_{0}:=\mu-\lambda\in\sigma(i\mathcal{K}_{P})$, so that there is at least one such $\omega$. But every other $\omega\in\sigma(i\mathcal{K})$ satisfies either $\omega>\omega_{0}$ or $\omega<\omega_{0}$. Therefore, for every other $\omega$, $\lambda+\omega>\mu$ or $\lambda+\omega<\mu$, and hence
@@
We consider a very simple 2 qubit system. Let $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ both be $\mathbbm{C}^{2}$ with the usual inner product. Let $|0\rangle:={\scriptscriptstyle\left(\begin{array}[]{c}0\cr 1\end{array}\right)}$ and $|1\rangle:={\scriptscriptstyle\left(\begin{array}[]{c}1\cr 0\end{array}\right)}$. Let $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ be the Pauli matrices, and define
@@ 8 Application to the computation of steady states
$P_{\mathcal{W}}=\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}e^{t\mathcal{K}_{P}}{\rm d}t\ .$
It follows that $P_{\mathcal{W}}$ is a CPTP operation, and hence $\|P_{\mathcal{W}}\|_{1\to 1}=1$.
@@
$\textnormal{Tr}_{A}[\mathcal{K}\bar{n}_{0}]=\textnormal{Tr}_{A}[-\mathcal{K}\mathcal{S}\mathcal{K}(\pi\otimes\bar{R})]+\textnormal{Tr}_{A}[\mathcal{K}\pi_{A}\otimes X_{0}]=\mathcal{D}_{P}\bar{R}+\mathcal{K}_{P}X_{0}=\mathcal{D}_{P}\bar{R}+\mathcal{K}_{P}V_{0}\ .$
Therefore, there is no constraint at this point on $W_{0}$, but $V_{0}$ must be chosen so that $\mathcal{K}_{P}V_{0}=-\mathcal{D}_{P}\bar{R}$, and this proves the first claim concerning the formula for $\bar{n}_{0}$ in (128).
Then with $\tilde{m}_{0}$ defined by (129), (128) becomes (130). By part (1) of Lemma 8.1, since $\bar{R}\in{\rm ker}(\mathcal{K}_{P})\cap{\rm ker}(\mathcal{D}_{P}^{\sharp})$, $\mathcal{D}_{P}(\bar{R})\in{\rm ran}(\mathcal{K}_{P})$, and then by part (2) of Lemma 8.1, there is a unique $V_{0}\in\mathcal{V}$ such that $\mathcal{K}V_{0}=-\mathcal{D}_{P}\bar{R}$, and $\|V_{0}\|_{1}\leqslant b^{-1}\|\mathcal{D}_{P}\bar{R}\|_{1}$, from which (131) follows. ∎
Now let $\bar{n}_{0}=\tilde{m}_{0}+\pi_{A}\otimes W_{0}$ as in (130) of Lemma 8.2, so that $\mathcal{D}\bar{n}_{1}=-\mathcal{K}\bar{n}_{0}$ is solvable. The general solution has the form
@@
Define $\tilde{m}_{1}:=-\mathcal{S}\mathcal{K}\bar{n}_{0}+\pi_{A}\otimes V_{1}=-\mathcal{S}\mathcal{K}\tilde{m}_{0}+\pi_{A}\otimes W_{0}+\pi_{A}\otimes V_{1}$}. For any choice of $W_{1}\in\mathcal{W}$, $\mathcal{D}\bar{n}_{2}=-\mathcal{K}(\tilde{m}_{1}+W_{1})$ is solvable, and
@@
For $1\leqslant\ell\leqslant k$, define $\tilde{m}_{\ell}:=-\mathcal{S}\mathcal{K}n_{\ell-1}+\pi_{A}\otimes V_{\ell}$, so that $\bar{n}_{\ell}=\tilde{m}_{\ell}+\pi_{A}\otimes W_{\ell}$, and let $\tilde{m}_{0}$ be given as in Lemma 8.2. Suppose that for all $1\leqslant\ell\leqslant k$,
@@
Then since $\|\bar{n}_{k}\|_{1}\leqslant\|\tilde{m}_{k}\|_{1}+\|W_{k}\|_{1}$,
@@
From (100) and (105), it follows that $\mathcal{D}_{P}(\pi_{A})=0$, and
$\mathcal{D}_{P}(\sigma_{1})=-\sigma_{1}\ ,\quad\mathcal{D}_{P}\sigma_{2}=-\sigma_{2}\quad{\rm and}\quad\mathcal{D}_{P}(\sigma_{3})=-2(\sigma_{3})\ ,$
with $t$ denoting $\tanh(\beta/2)$, $\mathcal{D}_{P}(\mathds{1})=-\tfrac{t}{\gamma}\sigma_{3}$. It follows that the matrix representation of the projected Lindblad generator $\mathcal{K}_{P}+\gamma^{-1}\mathcal{D}_{P}$ in the $\{\mathds{1},\sigma_{1},\sigma_{2},\sigma_{3}\}$ basis is
@@
Comparing (141) and (144), note that the real parts of the non-zero eigenvalues of $\gamma\mathcal{L}_{P,\gamma}$ are 1-1 and $\tfrac{3}{2}$, which are exactly the non-zero eigenvalues of $\mathcal{D}_{P}^{\sharp}$. This is consistent with, but does not imply, the relation between the mixing times in Theorem 7.5. Comparing (142) and (145), note that $\bar{R}=\lim_{\gamma\to\infty}\bar{R}_{\gamma}$.
@@ Appendix C Generalized inverses of Lindblad generators
$e^{t\mathcal{D}_{A}}=\sum_{j=0}^{n_{A}}e^{t\lambda_{j}}|Y_{j}\rangle\langle X_{j}|\ ,$
@@ Appendix D Volterra integral operators and the comparison of evolution equations
$\|\mathscr{H}\|_{\mathscr{C}\to\mathscr{C}}=\sup\{|\!|\!|\mathscr{H}A|\!|\!|\ :\ |\!|\!|A|\!|\!|\leqslant 1\}\ .$
@@
$\|\mathscr{H}-\widetilde{\mathscr{H}}\|_{\mathscr{C}\to\mathscr{C}}\leqslant\sup_{0\leqslant\tau\leqslant T}\left\{\ \int_{0}^{\tau}\|\mathcal{G}(\tau,\sigma)-\widetilde{\mathcal{G}}(\tau,\sigma)\|_{\mathcal{X}\to\mathcal{X}}{\rm d}\sigma\ \right\}\ .$
@@ Appendix E Mixing times
Let $\mathcal{L}$ be a Lindblad generator acting on $\widehat{\mathcal{H}}$, $\mathcal{H}$ a finite dimensional Hilbert space. By (163), and the definitions (24) and (25), if $\mathcal{L}$ is ergodic and gapped with unique steady state $\pi$, $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ for all $0<\epsilon<\tfrac{1}{2}$. The next lemma provides a converse.
@@
Let $\mathcal{L}$ be a Lindblad generator acting on $\widehat{\mathcal{H}}$. Assume only that $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ for some $0<\epsilon<\tfrac{1}{2}$. Then $\mathcal{L}$ has a unique steady state $\pi$, and ${\rm ker}(\mathcal{L})$ is spanned by $\pi$. Moreover, $\mathcal{L}$ has no purely imaginary eigenvalues so that in our finite dimensional setting $\mathcal{L}$ is ergodic and gapped. In particular, if $\mathcal{L}\rho=0$ and $\textnormal{Tr}[\rho]=1$, then $\rho=\pi$.
@@
Every Lindblad generator acting on $\widehat{\mathcal{H}}$ has at least one steady state, and since $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$, there is a unique steady state $\pi$. Since for all $X$, $\mathcal{L}(X^{\dagger})=\mathcal{L}(X)^{\dagger}$, if $X\in{\rm ker}(\mathcal{L})$ then $X^{\dagger}\in{\rm ker}(\mathcal{L})$. Hence if $\pi$ does not span ${\rm ker}(\mathcal{L})$ there would be a self-adjoint $X\in{\rm ker}(\mathcal{L})$ that is linearly independent of $\pi$. Subtracting an appropriate multiple of $\pi$ from $X$ and normalizing, we may assume that $\textnormal{Tr}[X]=0$ but $\|X\|_{1}=2$. Let $X=X_{+}-X_{-}$ be the spectral decomposition of $X$ into its positive and negative parts. Then $X_{+}$ and $X_{-}$ are both density matrices. Since $\mathcal{L}X=0$, for all $t>0$,
@@
Therefore, $2=\|X\|_{1}=\|e^{t\mathcal{L}}X_{+}-e^{t\mathcal{L}}X_{-}\|_{1}=2d_{{\rm TV}}(e^{t\mathcal{L}}X_{+},e^{t\mathcal{L}}X_{-})$. By (25), for $t>t_{{\rm mix}}(\mathcal{L},\epsilon)$, $d_{{\rm TV}}(e^{t\mathcal{L}}X_{+},e^{t\mathcal{L}}X_{-})<\epsilon$. This contradiction proves that $\pi$ span ${\rm ker}(\mathcal{L})$.
@@
We next show that when $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$, so that there is a unique stationary state $\pi$, $d_{{\rm TV}}(e^{kt_{{\rm mix}}(\mathcal{L})})$ decreases to zero exponentially fast in $k$. For $t>0$, let $P_{t}:=e^{t\mathcal{L}}$. In preparation for this, define the super-operator $\Pi$ by $\Pi X:=\textnormal{Tr}[X]\pi$. Then $\Pi$ is a CPTP map. It easy to write down an explicit Kraus representation showing this, but also $\Pi=\lim_{t\to\infty}P_{t}$, and each $P_{t}$ is CPTP. A super-operator $\mathcal{T}$ is said to be Hermitian in case $\mathcal{T}(X)^{\dagger}=\mathcal{T}(X^{\dagger})$ for all $X\in\widehat{\mathcal{H}}$. Since all CPTP operations are Hermitian (see e.g. [9, Lemma 5.2]), and hence $P_{t}-\Pi$ is Hermitian.
@@
Let $\rho_{0},\rho_{1}$ be any two density matrices on $\mathcal{H}$. For all $0<\epsilon<\tfrac{1}{2}$, and all $t\geqslant t_{{\rm mix}}(\mathcal{L},\epsilon)$, and all positive integers $k$,
@@
Now suppose that $t>t_{{\rm mix}}(\mathcal{L},\epsilon)$. By the definition (25) of $t_{{\rm mix}}(\mathcal{L},\epsilon)$, for any density matrix $\rho$,
@@
$t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ whenever $\mathcal{L}$ is ergodic and gapped. Moreover, by Theorem E.2 the definition of mixing time used here differs from the one defined in terms of (175) by at most a factor of $2$. Our applications require a formulation that does not refer to a putative unique invariant measure. Many variants of the mixing time are discussed in the classical literature on Markov chains, and the classical analog of the one considered here occurs in the foundational paper [2] of Aldous
@@
Data availability statement
In this work, we did not generate or analyze any datasets.
Conflict of interest statement
The authors declare that there are no competing interests with regard to this work.